There is a project with the following cash flows :

Year Cash Flow:

0 −$22,700

1 6,600

2 7,700

3 6,900

4 7,500

5 6,200

What is the payback period?

4 Years

3.75 Years

3.20 Years

3.45 Years

2.78 Years
Answer #1

Initial investment is 22700.

Till year 3 cashflow sum = 6600+7700+6900=20200 which is short of 22700 by 1500$

Year 4 cashflow = 7500

So it will take 1500/7500 of the 4th year for payback = 0.2

So payback = 3+0.2= 3.2 years